OpenStudy (anonymous):

Given the equation for simple harmonic motion d=2.3cos(4pi/3 t) a.find the maximum displacement b.find the frequency c.calculate the horizontal displacement at t=3 seconds

OpenStudy (anikhalder):

\[x=x_mcos(\omega t)\]\[x_m(\max.disp)=2.3m\]\[\omega =4 \pi/3=4.19 rads/\sec\]\[d(3 secs)=2.3\cos(\frac{ 4 \pi }{ 3}\times3)=2.3m\] Am I correct?

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (anikhalder):

Welcome